# TowerVision: Understanding and Improving Multilinguality in Vision-Language Models

### v2

@@ 1 Introduction
The success and widespread adoption of large language models (LLMs) has naturally led to a surge of interest in adding multimodal capabilities to these models. In particular, the visual modality has recently received considerable attention, with recent releases of frontier vision-language models (VLMs) (Deitke et al., 2024; OpenAI et al., 2024; Comanici et al., 2025; Team et al., 2025; Bai et al., 2025b). However, despite impressive progress, the development of VLMs has been mostly built upon English-centric language models, and trained with English vision-text data, giving little consideration to performance in most other languages. A key challenge in multilingualization of VLMs stems from an asymmetric data landscape—while high-quality text-only multilingual corpora are relatively abundant, high-quality multilingual vision-text data is scarce. As such, a critical challenge remains: What are the best strategies to effectively extend these models to support multiple languages beyond English?
An effective strategy for VLM multilingualization is to leverage large-scale text-only multilingual data to strengthen the language understanding of the text backbone, while complementing it with multimodal multilingual examples typically obtained through translation or high-quality synthetic generation, thereby reducing reliance on scarce real-world multilingual multimodal data. A recent example of this approach is Pangea (Yue et al., 2025), which extends multilinguality exclusively during the VLM fine-tuning stage using a mixture of data that combined multilingual vision-text pairs generated through synthetic data creation and machine translation of English instructions. While this strategy proved effective, it leaves open key questions: At which stages and on which modules should multilingualization be applied? Which design decisions yield the greatest impact? And how can visual grounding further enhance cross-lingual generalization?
@@
In this work, we introduce TowerVision,  [^1] The TowerVision model collection is available here. a suite of open-source multilingual VLMs built on top of Tower+ models (Rei et al., 2025) for 20 languages and dialects. [^2] English, German, Dutch, Spanish (Latin America), French, Portuguese (Portugal), Portuguese (Brazilian), Ukrainian, Hindi, Chinese (Simplified), Chinese (Traditional), Russian, Czech, Korean, Japanese, Italian, Polish, Romanian, Norwegian (Nynorsk) and Norwegian (Bokmål) To train TowerVision, we systematically address the challenges outlined above through comprehensive ablation studies, component-level analysis, and cross-lingual evaluation of a multilingualization recipe. Specifically, we investigate how to enhance the multilingual capabilities of VLMs from two axes: first, by performing several ablations to understand and explore the impact of the underlying components (including the alignment projector, vision encoder and text-only LLM); and second, by creating better, more multilingual vision-text datasets and exploring the impact of using this data across different VLM training stages. Overall, compared to strong VLMs of similar size, TowerVision exhibits competitive or superior performance on various multilingual and multimodal benchmarks, as well as cross-lingual transfer capabilities.
In addition to image-based VLMs, we also train a separate multilingual video model, TowerVideo, built on top of TowerVision, thereby extending our analysis to the video modality. TowerVideo achieves competitive performance on ViMUL-Bench (Shafique et al., 2025), a culturally-diverse multilingual video benchmark. Taken together, these contributions provide a comprehensive and systematic study of how to best integrate multilinguality into VLMs across modalities, architectural components, and training stages. Complementing the TowerVision family, we also release VisionBlocks  [^3] Released on Hugging Face upon acceptance., a curated dataset that consolidates and filters existing vision/video-language resources, further enriched with quality-controlled translations of English textual descriptions into 20 languages and dialects.
@@ 2 TowerVision
Our approach follows a multi-stage process encompassing three key components, illustrated in Figure 1: (i) a multilingual text-only backbone model, Tower+ Rei et al. (2025); (ii) a Vision Transformer encoder (ViT; Dosovitskiy et al. 2021) that processes visual inputs and extracts meaningful features; (iii) a connector/adapter module that transforms these visual features to generate representations compatible with the text embedding space. These modules can be selectively trained or kept frozen during different stages of development (Li et al., 2025). Although this training recipe and variations thereof are well-established and have produced several high-quality models (e.g., LLaVA (Liu et al., 2023b), Intern-VL (Chen et al., 2024), NVLM (Dai et al., 2024), Qwen2.5-VL (Bai et al., 2025b), Molmo (Deitke et al., 2024)), most of these fall short in capturing multilingual and culturally diverse nuances. We therefore introduce our multilingual adaptation, TowerVision—we first describe our carefully curated multilingual vision-text data, VisionBlocks2.1) and then describe the overall architecture and training procedure (§2.2).
@@ 2.1 VisionBlocks: Towards Better Multilingual Vision-Text Data
  • Human-written vision-text data featuring user-model interactions (common in text-only alignment) is severely limited. While abundant data exists from large-scale captioning datasets (e.g., LAION-5B; Schuhmann et al. 2022), such sources over prioritize scale over quality which is not ideal for training VLMs with advanced capabilities (Dong et al., 2025; Zhou et al., 2023) like instruction-following, helpfulness, and safety.
  • High-quality multilingual vision-text data is scarce; furthermore, the lack of open, high-quality multilingual VLMs makes controlled synthetic data challenging or restricted to closed models with limited usage licenses. The most viable alternative, also employed by Pangea (Yue et al., 2025), involves translating English vision-text interactions into target languages.
  • Filtering techniques such as reward model scoring or LLM-as-judge approaches (Gu et al., 2025) are significantly more challenging to implement for vision-text data, where even state-of-the-art VLMs (both open and proprietary) struggle to provide reliable preferences (Li et al., 2024).
@@
For English vision-text data, we use the mixture created in PixMo (Deitke et al., 2024) with a few minor changes: we exclude the AndroidControl, Points, and PointQA datasets, as they do not provide additional multilingual value at this stage; For multilingual vision-text data, we leverage a subset of “Open-Ended” and “Multiple-Choice” questions from CulturalGround (de Dieu Nyandwi et al., 2025) and the “Cultural” split of PangeaIns (Yue et al., 2025) for our languages of interest. The samples from PangeaIns are originally found in LAIONMulti (Schuhmann et al., 2022) that undergoes a series of automatic steps (using Gemini 1.5 Pro (Gemini Team et al., 2024)) including curating high-quality English instructions, carefully translating them to multiple languages, and adapting them for culturally-relevant multilingual contexts. CulturalGround uses a data curation pipeline that gathers culturally relevant entities from the Wikidata knowledge base, creates several questions and answers about each entity, rephrases them using an LLM, and filters low-quality samples using a VLM. In our work, we rely exclusively on CulturalGround’s filtered subsets to ensure maximum quality.
@@
In addition to the original English and multilingual captions, we translate the highly curated PixMo-Cap caption data Deitke et al. (2024) to our target languages using a Tower model (Alves et al., 2024). These translations are scored using CometKiwi (Rei et al., 2022) and filtered with a high threshold of 0.85 to ensure maximum quality. To further enhance diversity, we pair the remaining high-quality translations with a variety of language-specific captioning prompt templates (§A.5.1). We also augment the dataset with synthetic captions generated by the Gemini 2.5 API. For each image, we sample multiple system prompts to elicit diverse and detailed descriptions (see §A.5.2). This augmentation is intended to improve coverage of fine-grained visual details (e.g., spatial relations, attributes, and contextual cues) that human-authored captions often omit, and provides instruction-like supervision, aligning our model more closely with recent VLM training paradigms that leverage synthetic data to boost generalization and response quality. Similar strategies have been shown to be effective in scaling up instruction-following capabilities of VLMs such as LLaVA (Liu et al., 2023a) and InstructBLIP (Dai et al., 2023).
@@
To retain text-only performance of the backbone LLM, while acquiring strong multimodal capabilities through multimodal training requires a careful balance between text and multimodal data. Determining the optimal data mixture is non-trivial and typically involces a bunch of extensive ablation studies. For instance Pangea (Yue et al., 2025) include approximately 10% text-only data in their multimodal SFT mixture to retain text performance. In our work, we found that using around 20% text-only data provided the best trade-off between text-only and multimodal benchmark performance. We used EuroBlocks as the extended text-only data, a curated multilingual collection of high-quality synthetic data from the EuroLLM (Martins et al., 2025) synthetic post-training data. EuroBlocks provides diverse, instruction-aligned text that enriches our dataset with robust multilingual coverage and fine-grained, high-quality descriptions.
@@
As video-text data, we employ the LLaVA-Video-178k dataset (Zhang et al., 2025c), which contains captions alongside open-ended and multiple-choice questions in English. To make the dataset multilingual, we retain a randomly sampled half of the conversations in English, and we translate the remaining half uniformly into all supported languages using Tower+9B (Rei et al., 2025), thereby ensuring balanced cross-lingual coverage.
@@ 2.2 TowerVision: Architecture & Training Details
One way to improve the multilinguality of LLMs (e.g., improving cross-lingual understanding or extending multilingual support for other languages) is to start from a strong pretrained model and continue pretraining on carefully curated data, with subsequent post-training (Xu et al., 2024; 2025; Alves et al., 2024). TowerVision follows a similar principle, starting from a strong multilingual Gemma-based backbone Tower+ 2B/9B (Rei et al., 2025), which achieves strong multilingual general-purpose performance by leveraging a curated high-quality multilingual dataset and a training recipe designed to preserve general capabilities. As shown in §4, starting from this multilingual backbone substantially improves cross-lingual performance compared to starting from Gemma.
For the vision encoder, TowerVision is initialized with the recently proposed SigLIP2-so400m/14@384px (Tschannen et al., 2025), a vision transformer operating at $384\times 384$ resolution that extracts image patch representations and produces multilingually-aligned embeddings of size $729$. SigLIP2 is trained on a more diverse data mixture compared to alternatives such as CLIP-ViT (Radford et al., 2021), Perception Encoder (Bolya et al., 2025), or SigLIP1 (Zhai et al., 2023), and thereby yields better multilingual understanding, as we shall see in §4. To align the vision and text modalities, we use a LLaVA-based architecture (Liu et al., 2023b), where we train a projection layer consisting of a 2-layer MLP, randomly initialized. By combining Tower+ for text and SigLIP2 for vision, TowerVision benefits from complementary multilingual strengths across both modalities. The training process consists of three stages:
@@
    A vision finetuning phase, where we unfreeze the full model and train it on the full VisionBlocks dataset (§2.1), excluding the video-text data. In this phase, we use high-dynamic resolution (Liu et al., 2024a), breaking high-resolution images into a grid of smaller tiles which are then encoded with the vision encoder independently (together with a global thumbnail tile). The projected embeddings are then concatenated. We use a maximum of six tiles, which provides the best trade-off (§A.3). This phase leads to the TowerVision model.
@@
The models were trained on a custom fork of the LLaVA-Next (Liu et al., 2024a) codebase. [^4] https://github.com/deep-spin/LLaVA-NeXT
@@ 3.1 Tasks & Evaluation Benchmarks
We report results on ALM-Bench (Vayani et al., 2024), a cultural understanding multilingual [^5] German, Spanish, French, Italian, Korean, Dutch, Russian, English, Portuguese, Chinese (Simplified and Traditional), Icelandic, Czech, Ukrainian, Hindi, Japanese, Polish, Swedish, Hungarian, Romanian, Danish, Norwegian (Nynorsk), and Finnish. visual QA benchmark, OCRBench (Liu et al., 2024b) and cc-OCR (Yang et al., 2024) for English and multilingual [^6] German, French, Italian, Russian, Spanish, Korean, Portuguese. OCR-centric capabilities respectively, and TextVQA (Singh et al., 2019), assessing scientific understanding. Within cc-OCR, we report results on the multilingual text reading subset, as our primary focus is to evaluate the model’s multilingual text recognition capabilities.
@@
We report results on CoMMuTE (Futeral et al., 2023), a specialized multimodal translation benchmark that uses the visual content to resolve lexical ambiguities present in the source language, and Multi30K (Elliott et al., 2016), a standard benchmark for multimodal machine translation (MT) of image captions.
@@
We use ViMUL-Bench (Shafique et al., 2025), a multilingual video QA benchmark spanning 14 languages: Arabic (ar), Bengali (bn), Chinese (zh), English (en), French (fr), German (de), Hindi (hi), Japanese (ja), Russian (ru), Sinhala (si), Spanish (es), Swedish (sv), Tamil (ta), and Urdu (ur). The dataset contains both open-ended and multiple-choice questions covering culturally diverse domains such as festivals, customs, food, and heritage. Unlike prior datasets, ViMUL-Bench enables comprehensive evaluation of video-language models across both high- and low-resource languages, promoting inclusive and culturally aware research.

### v3

@@ 1 Introduction
The success and widespread adoption of large language models (LLMs) has naturally led to a surge of interest in adding multimodal capabilities to these models. In particular, the visual modality has recently received considerable attention, with recent releases of frontier vision-language models (VLMs) (deitke2024molmopixmoopenweights; openai2024gpt4ocard; comanici2025gemini; team2025gemma; bai2025qwen25vl). However, despite impressive progress, the development of VLMs has been mostly built upon English-centric language models, and trained with English vision-text data, giving little consideration to performance in most other languages. A key challenge in multilingualization of VLMs stems from an asymmetric data landscape—while high-quality text-only multilingual corpora are relatively abundant, high-quality multilingual vision-text data is scarce. As such, a critical challenge remains: What are the best strategies to effectively extend these models to support multiple languages beyond English?
An effective strategy for VLM multilingualization is to leverage large-scale text-only multilingual data to strengthen the language understanding of the text backbone, while complementing it with multimodal multilingual examples typically obtained through translation or high-quality synthetic generation, thereby reducing reliance on scarce real-world multilingual multimodal data. A recent example of this approach is Pangea (yue2025pangeafullyopenmultilingual), which extends multilinguality exclusively during the VLM fine-tuning stage using a mixture of data that combined multilingual vision-text pairs generated through synthetic data creation and machine translation of English instructions. While this strategy proved effective, it leaves open key questions: At which stages and on which modules should multilingualization be applied? Which design decisions yield the greatest impact? And how can visual grounding further enhance cross-lingual generalization?
@@
In this work, we introduce TowerVision,  [^1] The TowerVision model collection is available here. a suite of open-source multilingual VLMs built on top of Tower+ models (rei2025towerp) for 20 languages and dialects. [^2] English, German, Dutch, Spanish (Latin America), French, Portuguese (Portugal), Portuguese (Brazilian), Ukrainian, Hindi, Chinese (Simplified), Chinese (Traditional), Russian, Czech, Korean, Japanese, Italian, Polish, Romanian, Norwegian (Nynorsk) and Norwegian (Bokmål) To train TowerVision, we systematically address the challenges outlined above through comprehensive ablation studies, component-level analysis, and cross-lingual evaluation of a multilingualization recipe. Specifically, we investigate how to enhance the multilingual capabilities of VLMs from two axes: first, by performing several ablations to understand and explore the impact of the underlying components (including the alignment projector, vision encoder and text-only LLM); and second, by creating better, more multilingual vision-text datasets and exploring the impact of using this data across different VLM training stages. Overall, compared to strong VLMs of similar size, TowerVision exhibits competitive or superior performance on various multilingual and multimodal benchmarks, as well as cross-lingual transfer capabilities.
In addition to image-based VLMs, we also train a separate multilingual video model, TowerVideo, built on top of TowerVision, thereby extending our analysis to the video modality. TowerVideo achieves competitive performance on ViMUL-Bench (shafique2025culturallydiversemultilingualmultimodalvideo), a culturally-diverse multilingual video benchmark. Taken together, these contributions provide a comprehensive and systematic study of how to best integrate multilinguality into VLMs across modalities, architectural components, and training stages. Complementing the TowerVision family, we also release VisionBlocks  [^3] Released on Hugging Face upon acceptance., a curated dataset that consolidates and filters existing vision/video-language resources, further enriched with quality-controlled translations of English textual descriptions into 20 languages and dialects.
@@ 2 TowerVision
Our approach follows a multi-stage process encompassing three key components, illustrated in Figure 1: (i) a multilingual text-only backbone model, Tower+ rei2025towerp; (ii) a Vision Transformer encoder (ViT; dosovitskiy2021an) that processes visual inputs and extracts meaningful features; (iii) a connector/adapter module that transforms these visual features to generate representations compatible with the text embedding space. These modules can be selectively trained or kept frozen during different stages of development (li2025surveystateartlarge). Although this training recipe and variations thereof are well-established and have produced several high-quality models (e.g., LLaVA (liu2023visualinstructiontuning), Intern-VL (chen2024internvl), NVLM (dai2024nvlmopenfrontierclassmultimodal), Qwen2.5-VL (bai2025qwen25vl), Molmo (deitke2024molmopixmoopenweights)), most of these fall short in capturing multilingual and culturally diverse nuances. We therefore introduce our multilingual adaptation, TowerVision—we first describe our carefully curated multilingual vision-text data, VisionBlocks2.1) and then describe the overall architecture and training procedure (§2.2).
@@ 2.1 VisionBlocks: Towards Better Multilingual Vision-Text Data
  • Human-written vision-text data featuring user-model interactions (common in text-only alignment) is severely limited. While abundant data exists from large-scale captioning datasets (e.g., LAION-5B; schuhmann2022laion5bopenlargescaledataset), such sources over prioritize scale over quality which is not ideal for training VLMs with advanced capabilities (dong2025scalablevisionlanguagemodel; zhou2023limaalignment) like instruction-following, helpfulness, and safety.
  • High-quality multilingual vision-text data is scarce; furthermore, the lack of open, high-quality multilingual VLMs makes controlled synthetic data challenging or restricted to closed models with limited usage licenses. The most viable alternative, also employed by Pangea (yue2025pangeafullyopenmultilingual), involves translating English vision-text interactions into target languages.
  • Filtering techniques such as reward model scoring or LLM-as-judge approaches (gu2025surveyllmasajudge) are significantly more challenging to implement for vision-text data, where even state-of-the-art VLMs (both open and proprietary) struggle to provide reliable preferences (li2024vlrewardbenchchallengingbenchmarkvisionlanguage).
@@
For English vision-text data, we use the mixture created in PixMo (deitke2024molmopixmoopenweights) with a few minor changes: we exclude the AndroidControl, Points, and PointQA datasets, as they do not provide additional multilingual value at this stage; For multilingual vision-text data, we leverage a subset of “Open-Ended” and “Multiple-Choice” questions from CulturalGround (nyandwi2025groundingmultilingualmultimodalllms) and the “Cultural” split of PangeaIns (yue2025pangeafullyopenmultilingual) for our languages of interest. The samples from PangeaIns are originally found in LAIONMulti (schuhmann2022laion5bopenlargescaledataset) that undergoes a series of automatic steps (using Gemini 1.5 Pro (geminiteam2024gemini15unlockingmultimodal)) including curating high-quality English instructions, carefully translating them to multiple languages, and adapting them for culturally-relevant multilingual contexts. CulturalGround uses a data curation pipeline that gathers culturally relevant entities from the Wikidata knowledge base, creates several questions and answers about each entity, rephrases them using an LLM, and filters low-quality samples using a VLM. In our work, we rely exclusively on CulturalGround’s filtered subsets to ensure maximum quality.
@@
In addition to the original English and multilingual captions, we translate the highly curated PixMo-Cap caption data deitke2024molmopixmoopenweights to our target languages using a Tower model (alves2024tower). These translations are scored using CometKiwi (rei-etal-2022-cometkiwi) and filtered with a high threshold of 0.85 to ensure maximum quality. To further enhance diversity, we pair the remaining high-quality translations with a variety of language-specific captioning prompt templates (§A.5.1). We also augment the dataset with synthetic captions generated by the Gemini 2.5 API. For each image, we sample multiple system prompts to elicit diverse and detailed descriptions (see §A.5.2). This augmentation is intended to improve coverage of fine-grained visual details (e.g., spatial relations, attributes, and contextual cues) that human-authored captions often omit, and provides instruction-like supervision, aligning our model more closely with recent VLM training paradigms that leverage synthetic data to boost generalization and response quality. Similar strategies have been shown to be effective in scaling up instruction-following capabilities of VLMs such as LLaVA (liu2023llava) and InstructBLIP (dai2023instructblipgeneralpurposevisionlanguagemodels).
@@
To retain text-only performance of the backbone LLM, while acquiring strong multimodal capabilities through multimodal training requires a careful balance between text and multimodal data. Determining the optimal data mixture is non-trivial and typically involces a bunch of extensive ablation studies. For instance Pangea (yue2025pangeafullyopenmultilingual) include approximately 10% text-only data in their multimodal SFT mixture to retain text performance. In our work, we found that using around 20% text-only data provided the best trade-off between text-only and multimodal benchmark performance. We used EuroBlocks as the extended text-only data, a curated multilingual collection of high-quality synthetic data from the EuroLLM (martins2025eurollm) synthetic post-training data. EuroBlocks provides diverse, instruction-aligned text that enriches our dataset with robust multilingual coverage and fine-grained, high-quality descriptions.
@@
As video-text data, we employ the LLaVA-Video-178k dataset (zhang2025llavavideovideoinstructiontuning), which contains captions alongside open-ended and multiple-choice questions in English. To make the dataset multilingual, we retain a randomly sampled half of the conversations in English, and we translate the remaining half uniformly into all supported languages using Tower+9B (rei2025towerp), thereby ensuring balanced cross-lingual coverage.
@@ 2.2 TowerVision: Architecture & Training Details
One way to improve the multilinguality of LLMs (e.g., improving cross-lingual understanding or extending multilingual support for other languages) is to start from a strong pretrained model and continue pretraining on carefully curated data, with subsequent post-training (xu2024paradigm; xu2025xalma; alves2024tower). TowerVision follows a similar principle, starting from a strong multilingual Gemma-based backbone Tower+ 2B/9B (rei2025towerp), which achieves strong multilingual general-purpose performance by leveraging a curated high-quality multilingual dataset and a training recipe designed to preserve general capabilities. As shown in §4, starting from this multilingual backbone substantially improves cross-lingual performance compared to starting from Gemma.
For the vision encoder, TowerVision is initialized with the recently proposed SigLIP2-so400m/14@384px (tschannen2025siglip2multilingualvisionlanguage), a vision transformer operating at $384\times 384$ resolution that extracts image patch representations and produces multilingually-aligned embeddings of size $729$. SigLIP2 is trained on a more diverse data mixture compared to alternatives such as CLIP-ViT (radford2021learningtransferablevisualmodels), Perception Encoder (bolya2025perceptionencoderbestvisual), or SigLIP1 (zhai2023sigmoidlosslanguageimage), and thereby yields better multilingual understanding, as we shall see in §4. To align the vision and text modalities, we use a LLaVA-based architecture (liu2023visualinstructiontuning), where we train a projection layer consisting of a 2-layer MLP, randomly initialized. By combining Tower+ for text and SigLIP2 for vision, TowerVision benefits from complementary multilingual strengths across both modalities. The training process consists of three stages:
@@
    A vision finetuning phase, where we unfreeze the full model and train it on the full VisionBlocks dataset (§2.1), excluding the video-text data. In this phase, we use high-dynamic resolution (liu2024llavanext), breaking high-resolution images into a grid of smaller tiles which are then encoded with the vision encoder independently (together with a global thumbnail tile). The projected embeddings are then concatenated. We use a maximum of six tiles, which provides the best trade-off (§A.3). This phase leads to the TowerVision model.
@@
The models were trained on a custom fork of the LLaVA-Next (liu2024llavanext) codebase. [^4] https://github.com/deep-spin/LLaVA-NeXT
@@ 3.1 Tasks & Evaluation Benchmarks
We report results on ALM-Bench (vayani2024alm), a cultural understanding multilingual [^5] German, Spanish, French, Italian, Korean, Dutch, Russian, English, Portuguese, Chinese (Simplified and Traditional), Icelandic, Czech, Ukrainian, Hindi, Japanese, Polish, Swedish, Hungarian, Romanian, Danish, Norwegian (Nynorsk), and Finnish. visual QA benchmark, OCRBench (Liu_2024) and cc-OCR (yang2024ccocrcomprehensivechallengingocr) for English and multilingual [^6] German, French, Italian, Russian, Spanish, Korean, Portuguese. OCR-centric capabilities respectively, and TextVQA (singh2019text_vqa), assessing scientific understanding. Within cc-OCR, we report results on the multilingual text reading subset, as our primary focus is to evaluate the model’s multilingual text recognition capabilities.
@@
We report results on CoMMuTE (futeral-etal-2023-tackling), a specialized multimodal translation benchmark that uses the visual content to resolve lexical ambiguities present in the source language, and Multi30K (elliott-etal-2016-multi30k), a standard benchmark for multimodal machine translation (MT) of image captions.
@@
We use ViMUL-Bench (shafique2025culturallydiversemultilingualmultimodalvideo), a multilingual video QA benchmark spanning 14 languages: Arabic (ar), Bengali (bn), Chinese (zh), English (en), French (fr), German (de), Hindi (hi), Japanese (ja), Russian (ru), Sinhala (si), Spanish (es), Swedish (sv), Tamil (ta), and Urdu (ur). The dataset contains both open-ended and multiple-choice questions covering culturally diverse domains such as festivals, customs, food, and heritage. Unlike prior datasets, ViMUL-Bench enables comprehensive evaluation of video-language models across both high- and low-resource languages, promoting inclusive and culturally aware research.